OpenStudy (anonymous):

PLEASE HELP IM SO CONFUSED write an equation to solve each problem 1.)two buses leave Houston at the same time and travel opposite directions. one bus averages 55mi/h and the other bus averages 45mi/h. when well they be 400 mi apart? 2.) two planes left an airport at noon. one flew east and the other flew west at twice the speed. After 3 hours the planes were 2700 miles apart. how fast was each plane flying?

OpenStudy (abb0t):

Divide 400 by 55. That will be your total hours. Use the units to cancel out to solve for what you want, hrs. So you are trying to solve for hours, given miles, \(400\frac{mi}{1} \times \frac{h}{55~mi}\) So you notice that the miles cancel out! And you re left with \(\frac{400}{55} h\) Upon dividing, you get 4 hours.

OpenStudy (anonymous):

What? when you divide 400 by 55 you get 7.272...

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

Oh i get it!!! you add the 45 and 55 than divide!

OpenStudy (anonymous):

because they are two different buses and when you add they equal 100 than you take 400/100 and get 4 hours like what you said...

OpenStudy (abb0t):

Well, I just divide each one individually, 400 divided by 55 = 4

OpenStudy (jdoe0001):

\(\begin{matrix} &distance& rate & time\\ 1st\ bus& x & 55 & t\\ \hline\\ 2nd\ bus&400-x&45&t \end{matrix}\)

OpenStudy (anonymous):

400 divided by 55 = 7.272727273, check your calculator again

OpenStudy (jdoe0001):

see the equations?

OpenStudy (jdoe0001):

when they both reach 400 miles apart, their time is the same, thus the time for 1st bus = time for 2nd bus

OpenStudy (abb0t):

oh, wait, you're right. Lol. i did this in my head. LOL. i was getting 200 for some reason.

OpenStudy (jdoe0001):

how far has 1st bus traveled? we dunno, thus "x" how far has 2nd bus traveled? well, if trhe 1st bust traveled "x", the 2nd one traveled 400-x

OpenStudy (anonymous):

ok so how do we find x

OpenStudy (anonymous):

how would i make one equation with that info

OpenStudy (jdoe0001):

\(\begin{matrix} &distance& rate & time\\ 1st\ bus& x & 55 & t\\ \hline\\ 2nd\ bus&400-x&45&t\\ \hline\\ \hline\\ \textit{distance = rate } \times \textit{time}\\ \bf x = 55t\\ \bf 400-x = 45t \end{matrix}\)

OpenStudy (jdoe0001):

notice that x= 55t on 1st of the system of equations since we know that now then, use that value of "x" in the 2nd equation, what do you get?

OpenStudy (anonymous):

you would add 55t on both sides and get 100t than divide by 100 on both sides and get t= 4 hours

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

thank you so much! i can do this for the second one too :)

OpenStudy (jdoe0001):

yeah, the 2nd is pretty much the same thing

OpenStudy (anonymous):

with the second one would the equation be 2700=3x+2x(3) or???

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

yay!! i got it thank you :)

OpenStudy (jdoe0001):

yw